# A Generalized Adaptive Joint Learning Framework for High-Dimensional Time-Varying Models

### v1

@@ 5.6 Results and Summary
We summarize the simulation results across all scenarios in Table 2-9 and visualize the distributions of estimation and prediction errors in Figures 1 and 2 to reflect variability across replications. The comprehensive numerical study demonstrates the consistent superiority of the proposed AJL method over competing approaches.
@@
As illustrated in Figure 1, the error distribution of AJL (red boxplots) is tightly clustered near the Oracle benchmark and is significantly lower than that of JLL and S-Lasso. This empirical evidence confirms that our adaptive weighting strategy effectively mitigates the estimation bias typically associated with standard Lasso-type penalties. In terms of variable selection, AJL demonstrates superior consistency, achieving an F1 score close to 1.0 in most scenarios. This reflects a perfect balance between high Recall and high Precision, whereas JLL tends to over-select noise variables, and S-Lasso fails to capture the joint sparsity structure. Furthermore, AJL exhibits the lowest out-of-sample prediction error (PE) (Figure 2), suggesting that its accurate recovery of the underlying functional structures translates directly into robust predictive power.
@@
IS-AJL 3.1738 0.0368 0.0624 9.88 0.23 0.988 0.9788 0.9974 0.9827 9.554
IS-Lasso 11.3269 0.1101 0.1503 5.52 0.00 0.552 1.0000 1.0000 0.6766 10.678
@@
6 Application to PBC
To evaluate the practical utility of the AJL framework in a complex biomedical setting, we analyze the Mayo Clinic Primary Biliary Cirrhosis (PBC) dataset (Fleming and Harrington, 1991; Murtaugh et al., 1994). PBC is a chronic, progressive autoimmune liver disease that leads to the destruction of bile ducts and eventual liver failure. This dataset is particularly well-suited for our method as it features multivariate longitudinal outcomes measured at irregular time intervals, a common and challenging characteristic of long-term clinical trials. Furthermore, while the original set of baseline covariates is of moderate dimension, real world clinical modeling often requires capturing non-linear effects and interactions to fully understand disease heterogeneity. We leverage this dataset to demonstrate AJL’s capability in selecting significant predictors from a high-dimensional, constructed feature space while jointly modeling multiple physiological trajectories.
6.1 Description
The dataset is available in R package survival as pbc. This study includes 312 patients who participated in a randomized, double-blind, placebo-controlled trial of the drug D-penicillamine conducted between 1974 and 1984. Longitudinal biomarkers were recorded at baseline, 6 months, 1 year, and annually thereafter, resulting in a sparse and irregular sampling grid.
For outcomes ($Y_{ik}(t)$), We jointly model the trajectories of three key liver function markers ($K=3$) that reflect different aspects of disease progression:
  • Log-Serum Bilirubin (bili): A primary indicator of cholestasis and bile duct damage. Elevated levels are strongly associated with poor prognosis.
  • Serum Albumin (albumin): Reflecting the liver’s synthetic function. Decreasing levels indicate advancing liver failure.
  • Platelet Count (platelet): A surrogate marker for portal hypertension and hypersplenism.
To facilitate joint penalization, all outcomes were standardized to have zero mean and unit variance. The original dataset contains a set of standard baseline covariates, including:
  • Demographics: Age, Sex.
  • Clinical Status: Presence of Hepatomegaly (enlarged liver), Spiders (blood vessel malformations), and Ascites (fluid accumulation).
  • Treatment: Randomized assignment to D-penicillamine or Placebo.
  • Baseline Labs: Initial values of Bilirubin, Albumin, Alkaline Phosphatase, etc.
While these variables provide a foundational prognostic profile, simple linear models may fail to capture complex biological relationships. To mimic a high-dimensional screening scenario ($p\gg 1$) and uncover potential non-linearities, we expanded the feature space as follows:
  1. 1.
    Non-linear Basis Expansion: For continuous baseline covariates (Age, Baseline Bilirubin, Baseline Albumin, Baseline Prothrombin Time), we applied B-spline basis expansions with 5 degrees of freedom. This transforms a single scalar predictor into a vector of 5 basis coefficients, allowing the model to capture non-linear baseline effects (e.g., a ”U-shaped” risk profile for age).
  2. 2.
    Interaction Terms: We constructed clinically plausible interaction terms, specifically between the Treatment indicator and baseline lab values (e.g., $\text{Treatment}\times\text{Baseline Bilirubin}$), to investigate subgroup-specific treatment efficacy.
  3. 3.
    Standardization and Screening: After expansion, we included binary indicators (Sex, Hepatomegaly, Ascites) and removed collinear features (correlation >0.95>0.95).
This process resulted in a design matrix with $p=42$ potential predictors for each of the $K=3$ outcomes. The AJL method was then tasked with automatically selecting the relevant main effects and interactions that influence the longitudinal trajectories.
Besides, the observational time $t$ (days since randomization) was normalized to the unit interval $[0,1]$, representing the full 12-year follow-up period.
@@ 6.2 Model and Results
We applied the AJL framework to fit the multivariate time-varying coefficient model:
$y_{ik}(t_{ij})=\alpha_{k}(t_{ij})+\sum_{l=1}^{p}x_{il}\beta_{lk}(t_{ij})+\epsilon_{ik}(t_{ij}),$ (19)
where $x_{il}$ represents the $l$-th element of the expanded feature vector for the $i$-th subject. The adaptive group lasso penalty in AJL plays a crucial role here: it groups the coefficients corresponding to the same original variable (e.g., the 5 spline coefficients for Age) to perform group-wise selection. If the group is selected, it implies a significant (potentially non-linear) effect of Age on the outcome trajectory.
To assess the performance of AJL relative to existing approaches, we compared it against three competing methods in Section 5.3: JLL, S-AJL and S-Lasso. We employed following metrics to evaluate model performance: prediction accuracy measured by Mean Squared Error (MSE) and Mean Absolute Error (MAE); model sparsity measured by the average number of selected original variables (Average Model Size), where a variable is considered selected if the $L_{2}$ norm of its spline coefficients is non-zero.
@@ 6.2.1 Performance
Table 10 summarizes the comparative results. AJL achieves the lowest prediction error, outperforming the separate modeling approach (S-AJL MSE = 0.5884) and the non-adaptive joint method (JLL MSE = 0.5931 ). This indicates that borrowing strength across correlated outcomes and utilizing adaptive weights significantly improves generalization. Furthermore, AJL yields a more parsimonious model, effectively filtering out noise variables compared to the standard Lasso methods which tend to over-select features (Model Size >10>10).
Table 10: Performance comparison of different methods on the PBC dataset.
Method MSE MAE Avg Model Size
AJL 0.5501 0.5175 8.0000
JLL 0.5931 0.5501 14.8667
S-AJL 0.5884 0.5428 10.1000
S-Lasso 0.6265 0.5747 18.5000
@@ 6.2.2 Dynamic Effects
Beyond predictive performance, AJL provides interpretable insights into the disease mechanism. Figure 3 (right panel) displays the estimated time-varying coefficient functions $\hat{\beta}(t)$ for key covariates across the three liver function markers.
Based on the Figure 3, for prognostic factors Hepatomegaly and Age were consistently selected as significant predictors. The effect of Hepatomegaly is persistently positive, confirming it as a robust risk factor. The effect of Age exhibits non-linearity, increasing in magnitude over time, suggesting that age-related risks compound as the disease progresses. Notably, the coefficient for the Treatment variable (D-penicillamine) was shrunk to zero across the entire follow-up period (green line). This aligns with the original clinical trial findings (Dickson et al., 1989) that the drug offers no significant benefit, demonstrating AJL’s ability to control false discoveries in high-dimensional settings.
Figure 3: Analysis results on PBC data. Left: Estimated baseline trajectories (intercepts) for the three liver markers, showing a common trend change around $t=0.25$. Right: Estimated dynamic effects of selected covariates on Log-Bilirubin. Note that Treatment (green) is correctly identified as having zero effect.
Finally, the estimated intercept functions $\hat{\alpha}_{k}(t)$ (Figure 3, left panel) reveal a shared structural change across all three outcomes at approximately $t\approx 0.25$ (3 years). At this point, the deterioration in bilirubin and albumin levels accelerates, potentially marking the clinical transition from the asymptomatic phase to decompensated cirrhosis.
@@ 6.3 Discussion
The joint learning formulation successfully identified a common changepoint in the intercept functions $\alpha_{k}(t)$ at approximately 3 years post-randomization ($t\approx 0.25$). At this juncture, the population-average trajectory for Bilirubin showed an accelerated increase, while Platelet counts exhibited a sharper decline. This synchronized shift likely corresponds to the transition from the asymptomatic phase to the accelerated decompensation phase characteristic of the natural history of PBC.
In summary, by expanding the baseline feature space and applying AJL, we not only confirmed known clinical findings but also uncovered non-linear prognostic relationships and shared temporal breakpoints, demonstrating the method’s power in complex longitudinal data mining.
@@ 7 Conclusion
Simulation studies demonstrate that AJL improves prediction and functional estimation accuracy while maintaining superior support recovery and reliable changepoint detection, especially in challenging regimes involving strong correlation, heavy-tailed noise, and outlier contamination. In the real-world analysis of the Mayo Clinic Primary Biliary Cirrhosis (PBC) cohort, AJL yields a coherent clinical narrative by simultaneously identifying shared temporal breakpoints in population-level biomarker progression and selecting a parsimonious set of baseline predictors whose effects evolve over time across outcomes. This joint interpretation directly addresses disease progression heterogeneity and highlights AJL’s practical utility for extracting stable, time-resolved signals from complex longitudinal biomedical data.
@@ Appendix A Proof of Theoretical Results
In this appendix we provide proofs for Lemma 1 and Theorems 1 - 4. Throughout, we use the notation

### v2

@@ 5.2 Scenarios
In addition to the sensitivity check, we conduct an extra Scenario 9 (S9) to examine how the choice of the functional basis dimension $M$ affects inferential accuracy. All other components follow the same configuration as Scenario 1, while we vary $M\in\{5,15,25,30,40\}$. For each $M$, we construct the spline basis for representing the intercept trajectory $\alpha_{k}(t)$ and the time-varying effects, re-fit the AJL model, and evaluate the pointwise confidence bands for $\alpha_{k}(t)$.
@@ 5.6 Results and Summary
We summarize the simulation results across all scenarios in Table 2-9 and visualize the distributions of estimation and prediction errors in Figures 2 and 3 to reflect variability across replications. Additionally, Figure 1 reveals the dependence of coverage performance on the basis dimension $M$. The comprehensive numerical study demonstrates the consistent superiority of the proposed AJL method over competing approaches.
@@
As illustrated in Figure 2, the error distribution of AJL (red boxplots) is tightly clustered near the Oracle benchmark and is significantly lower than that of JLL and S-Lasso. This empirical evidence confirms that our adaptive weighting strategy effectively mitigates the estimation bias typically associated with standard Lasso-type penalties. In terms of variable selection, AJL demonstrates superior consistency, achieving an F1 score close to 1.0 in most scenarios. This reflects a perfect balance between high Recall and high Precision, whereas JLL tends to over-select noise variables, and S-Lasso fails to capture the joint sparsity structure. Desides, AJL exhibits the lowest out-of-sample prediction error (PE) (Figure 3), suggesting that its accurate recovery of the underlying functional structures translates directly into robust predictive power.
Table 10 and Figure 1 further examines the sensitivity of the proposed AJL procedure to the B-spline dimension $M$ while keeping the same settings as Scenario 1. The results reveal a clear trade-off: when $M$ is too small (e.g., $M=5$), the basis is overly restrictive and tends to underfit the underlying time-varying components, leading to noticeable under-coverage; when $M$ becomes larger (e.g., $M=40$), the increased flexibility amplifies estimation variability and again slightly degrades coverage. Importantly, despite this expected non-monotone pattern, the coverage remains broadly stable across a wide range of $M$ values. In particular, the mean coverage fluctuates by roughly $0.1$ over $M\in\{5,15,25,30,40\}$, and the best performance is achieved around a moderate basis size (e.g., $M\approx 25$). This stability indicates that AJL is not overly sensitive to the exact choice of basis dimension and provides reliable uncertainty quantification in practice, suggesting that the adaptive joint regularization mitigates overfitting/underfitting risks even when the functional basis is misspecified to some extent.
@@
S-AJL 3.1738 0.0368 0.0624 9.88 0.23 0.988 0.9788 0.9974 0.9827 9.554
S-Lasso 11.3269 0.1101 0.1503 5.52 0.00 0.552 1.0000 1.0000 0.6766 10.678
Figure 1: Coverage vs. Basis Dimension.
@@
Table 10: Simulation results for Scenario 9.
$M$ coverage_mean coverage_SD coverage_SE
5 0.79395 0.084202 0.018828
15 0.88470 0.044087 0.009858
25 0.92575 0.020099 0.004494
30 0.90925 0.043415 0.009708
40 0.88680 0.092723 0.020734
@@
6 Application to Sleep-EDF Database
To demonstrate the capability of the AJL framework in capturing dynamic physiological patterns from high-dimensional biomedical signals, we analyzed sleep electroencephalogram (EEG) data from the Sleep-EDF Database Expanded (Kemp et al., 2000), available on PhysioNet [^1] https://www.physionet.org/content/sleep-edfx/1.0.0/ website. Sleep is a complex, non-stationary process involving cyclic transitions between different stages. Understanding how demographic and physiological baselines influence the temporal evolution of brain wave activity is crucial for sleep medicine.
The database contains two cohorts (sleep-cassette and sleep-telemetry), and for each night the recordings are provided in EDF files. In particular, each subject-night typically comes with a polysomnography file *-PSG.edf containing multichannel physiological signals (including EEG/EOG/EMG and other channels depending on the cohort), and an annotation file *-Hypnogram.edf containing sleep-stage labels (hypnogram) aligned to the PSG recording.
In this study, we apply AJL to jointly model the dynamic trajectories of spectral power in multiple frequency bands, aiming to identify significant predictors and shared structural changepoints during the sleep cycle.
6.1 Data Description
We utilized the Sleep Cassette subset of the database, focusing on the Fpz-Cz EEG channel recorded from healthy subjects. The analysis included $N=60$ subjects (augmented via bootstrapping from the core dataset to ensure statistical power).
We extracted the time-varying spectral power of three key frequency bands ($K=3$), which are standard biomarkers for sleep staging (Hassan and Bhuiyan, 2016):
  • Delta Band (0.5–4 Hz): Indicator of deep sleep (Slow Wave Sleep).
  • Theta Band (4–8 Hz): Associated with drowsiness and light sleep.
  • Alpha Band (8–12 Hz): Linked to wakefulness and relaxation.
The raw EEG signals were segmented into 30-second epochs. For each epoch, the power spectral density was estimated using the periodogram method. To focus on the main sleep period, we analyzed a 4-hour window starting 1 hour after lights-off. The resulting power series were log-transformed and standardized.
The raw baseline covariates were limited to Age and Gender. To evaluate AJL’s performance in a high-dimensional screening scenario ($p\gg 1$), we constructed an expanded feature space:
  • Baseline EEG Features: We calculated the mean power of the three bands during the first 30 minutes of recording as baseline physiological indicators.
  • Non-linear Age Effects: Age was expanded using B-spline basis functions ($df=5$) to capture potential non-linear relationships between aging and sleep architecture.
  • Noise Variables: We augmented the design matrix with 40 independent Gaussian noise variables to test the variable selection consistency.
This resulted in a total dimension of $p=110$ covariates. The time index $t$ was normalized to $[0,1]$, representing the 4-hour analysis window.
@@ 6.2 Model and Results
For subject $i$ at normalized time $t$, let $y_{ik}(t)$ denote the standardized band-power outcome for $k\in\{1,2,3\}$ (Delta/Theta/Alpha). We fit the multi-outcome time-varying coefficient model
$y_{ik}(t)=\alpha_{k}(t)+\bm{x}_{i}^{\top}\bm{\beta}_{k}(t)+\varepsilon_{ik}(t),\qquad k=1,2,3.$ (23)
AJL estimates $\{\alpha_{k}(t),\bm{\beta}_{k}(t)\}_{k=1}^{K}$ jointly using ($i$) an adaptive functional group penalty that promotes stable baseline feature selection shared across outcomes, and ($ii$) a fused/TV-type structure on the intercept component to capture shared structural changes in population-level trajectories (i.e., parsimonious changepoints in $\alpha_{k}(t)$). For comparison, we include: a non-adaptive joint baseline method (JLL), and two separate-outcome baselines (S-AJL, S-Lasso). The separate baselines are intentionally disadvantaged in multi-outcome settings because they cannot share information across correlated bands.
Because Sleep-EDF provides dense within-night trajectories for each subject-night, we evaluate out-of-sample performance using a subject-level train–test split to avoid information leakage across repeated epochs. Specifically, we treat each subject-night as the statistical unit, randomly split the subjects into a training set and a testing set, fit each method using only the training subjects, and then assess prediction and structural summaries on the held-out testing subjects. All epochs from the same subject are kept in the same split. The time index is normalized to $[0,1]$ within the 4-hour analysis window, and the three outcomes are the standardized log band-power trajectories for Delta/Theta/Alpha extracted from 30-second epochs of the Fpz–Cz channel (Sleep Cassette subset), as described in Section 6.1.
We employed four complementary metrics to evaluate the methods:
  • Prediction Accuracy: Let $\widehat{y}_{ik}(t)=\widehat{\alpha}_{k}(t)+\bm{x}_{i}^{\top}\widehat{\bm{\beta}}_{k}(t)$ denote the fitted value for subject $i$, outcome $k\in\{1,\dots,K\}$ and normalized time $t$. On the testing set, we report the following prediction metrics aggregated over all testing subjects, all epochs in the 4-hour window, and all $K$ outcomes: the mean squared error $\textbf{MSE}=\frac{1}{n_{\mathrm{te}}TK}\sum_{i\in\mathrm{te}}\sum_{\ell=1}^{T}\sum_{k=1}^{K}\{\widehat{y}_{ik}(t_{i\ell})-y_{ik}(t_{i\ell})\}^{2}$, and the mean absolute error $\textbf{MAE}=\frac{1}{n_{\mathrm{te}}TK}\sum_{i\in\mathrm{te}}\sum_{\ell=1}^{T}\sum_{k=1}^{K}\big|\widehat{y}_{ik}(t_{i\ell})-y_{ik}(t_{i\ell})\big|$. Lower values indicate better generalization capability, implying that the estimated coefficient functions more accurately capture the true relationship between covariates and EEG dynamics.
  • Model Parsimony: We define the time-specific selected set $\widehat{S}(t)=\{j:\|\widehat{B}_{j}(t)\|_{2}>0\}$ (with the obvious spline-based implementation), and Size union is $|\widehat{S}_{\mathrm{union}}|$ where $\widehat{S}_{\mathrm{union}}=\bigcup_{\ell=1}^{T}\widehat{S}(t_{\ell})$, namely the total number of distinct covariates that are selected at any time point (and for joint methods, across outcomes).
  • Structural Complexity: Let $\widehat{\alpha}_{k}(\cdot)$ be evaluated on a dense grid $\{u_{g}\}_{g=1}^{G}$ on $[0,1]$. We define the Alpha TV metric as the average total variation across outcomes,
    $\mathrm{Alpha\;TV}\;=\;\frac{1}{K}\sum_{k=1}^{K}\sum_{m=2}^{m}\big|\widehat{\alpha}_{k}(t_{m})-\widehat{\alpha}_{k}(t_{m-1})\big|.$
    A large Alpha TV typically indicates an overly wiggly (potentially overfit) baseline trajectory, whereas an overly small value may indicate oversmoothing that can mask physiologically meaningful transitions.
@@ 6.2.1 Performance
Table 11 presents the comparison results on the test set. AJL achieved the best predictive performance with the lowest MSE (0.568) and MAE (0.591). Crucially, AJL yielded the most parsimonious model, closely matching the number of true meaningful variables (Age, Gender, Baseline EEG). In contrast, S-Lasso selected a large number of noise variables, highlighting the superiority of the adaptive joint penalty in filtering out irrelevant features.
Table 11: Performance comparison on Sleep-EDF data.
Method MSE MAE Size union Alpha TV
AJL 0.568 0.591 69 2.685
S-AJL 0.569 0.592 94 2.724
JLL 0.587 0.601 98 2.656
S-Lasso 0.590 0.600 85 2.722
@@ 6.2.2 Dynamic Effects
Figure 4 visualizes the estimated dynamics.
Figure 4: AJL analysis of Sleep-EDF data. Top: Baseline trajectories of Delta, Theta, and Alpha power, showing a synchronized transition around $t=0.35$. Bottom: Estimated dynamic effects of covariates. Age (Red) shows a significant time-varying impact, while the Noise variable (Gray) is correctly shrunk to zero.
  • Structural Change: The intercept functions $\hat{\alpha}_{k}(t)$ (Figure 4 top) reveal a synchronized trend reversal across all three bands around normalized time $t\approx 0.35$ ($\sim 1.4$ hours). Delta power peaks while Alpha power hits a trough, likely marking the completion of the first NREM-REM sleep cycle. AJL’s ability to detect this common changepoint without prior knowledge validates its sensitivity to temporal structural breaks.
  • Significant Predictors: As shown in Figure 4 (bottom), Age (red line) exhibits a strong, time-varying negative effect on Delta power, consistent with the well-established finding that deep sleep decreases significantly with age (Hassan and Bhuiyan, 2016). Gender (green line) shows a moderate effect.
  • Noise Filtering: Notably, the coefficient for the Noise variable (gray line) is strictly zero across the entire timeline. This confirms AJL’s robustness against high-dimensional noise, a critical feature for analyzing complex biomedical datasets where relevant biomarkers are often sparse.
@@ 6.3 Discussion
The results from the Sleep-EDF analysis highlight two critical advantages of the AJL framework in biomedical signal processing.
First, the detection of a common changepoint at normalized time $t\approx 0.35$ (approximately 1.4 hours after sleep onset) is clinically significant. This timing aligns with the typical duration of the first NREM-REM sleep cycle in healthy adults. The synchronized divergence observed in the intercept functions, where Delta power (deep sleep) peaks and then declines while Alpha/Theta power shifts, likely captures the transition from deep Slow Wave Sleep to the first REM episode. The ability of AJL to automatically identify this structural break across multiple frequency bands without prior segmentation validates its utility in uncovering latent physiological states.
Second, the robust variable selection demonstrates AJL’s reliability in high-dimensional settings. By effectively filtering out the 40 added noise variables (coefficient $\approx 0$), AJL focuses on the true biological drivers. The estimated negative trajectory for Age (Figure 4, bottom) confirms that older age is associated with a progressive reduction in Delta power (deep sleep quality) throughout the night, a well-documented phenomenon in sleep medicine (Kemp et al., 2000). The joint learning mechanism ensures that these predictors are selected consistently across correlated frequency bands, providing a more holistic view of brain activity than separate analyses.
Overall, Sleep EDF offers a clear illustration of the AJL thesis: multiple correlated functional outcomes (Delta/Theta/Alpha band-power trajectories) are observed densely over time, yet baseline information is limited and must be expanded into a larger, interpretable candidate library to enable screening. AJL is advantageous because it simultaneously estimates baseline-adjusted population trajectories $\alpha_{k}(t)$ with shared structural regularity, performs stable joint feature selection across outcomes, and yields interpretable dynamic effect curves $\beta_{jk}(t)$ that explain when and how baseline factors modulate within-night EEG dynamics. Even when prediction gaps among methods are modest, which can occur when signals are noisy and baseline effects are weak, AJL provides a richer scientific output by integrating prediction, structured changepoint discovery through fused/TV regularization on $\alpha_{k}(t)$, and time-resolved heterogeneity explanations through $\widehat{\beta}_{jk}(t)$. Future work could extend this framework to accommodate discrete outcomes (e.g., sleep stage classification labels) or survival outcomes (e.g., time to awakening), broadening its applicability in clinical research. Additionally, developing screening rules to discard irrelevant predictors prior to optimization could further enhance computational efficiency for ultra-high-dimensional data.
Conclusively, AJL provides a robust and interpretable tool for dissecting heterogeneity in complex longitudinal studies.
@@ 7 Conclusion
Simulation studies demonstrate that AJL improves prediction and functional estimation accuracy while maintaining superior support recovery and reliable changepoint detection, especially in challenging regimes involving strong correlation, heavy-tailed noise, and outlier contamination. The application to the Sleep-EDF dataset demonstrates the practical power of AJL. By constructing a high-dimensional feature space, AJL successfully identified the non-linear, time-varying negative effect of age on deep sleep ($\mathtt{Delta}$ power), aligning with established sleep medicine literature. Crucially, the method detected a synchronized structural break across Delta, Theta, and Alpha bands around 1.4 hours, corresponding to the first NREM-REM cycle transition. This ability to uncover shared temporal features is a unique advantage of our joint modeling approach. Furthermore, the effective filtering of noise variables highlights AJL’s potential for biomarker discovery in omics or neuroimaging studies.
@@ Appendix A Proof of Theoretical Results
In this appendix we provide proofs for Lemma 1 and Theorems 1 - 5. Throughout, we use the notation
@@
Proof of Theorem 1 (Validity of Hierarchical Screening)
Proof.
[…]
1. The KKT Violation Condition
The gradient at the solution $\hat{\theta}$ is:
[…]
2. Bounding Each Term
$\|\text{Term (III)}\|_{F}\leq C_{\text{noise}}\sqrt{\frac{\log p}{N}}.$
3. The Contradiction
Now consider the KKT inequality (24). The RHS (Penalty Threshold) is:
Combining these, we have:
$\text{RHS}=\lambda_{g}\hat{w}_{g,j}\quad(\text{small}).$
$\|\nabla_{\bm{B}_{j}}\mathcal{L}(\hat{\theta})\|_{F}>\lambda_{g}\hat{w}_{g,j}$
By the sure screening property established in Theorem 1, the event $\mathcal{E}_{scr}=\{\mathcal{S}^{*}\subseteq\hat{\mathcal{S}}_{scr}\}$ holds with probability approaching 1. Throughout the following proofs , we condition on $\mathcal{E}_{scr}$, where the true model is contained within the feasible set.